OpenStudy (anonymous):

Find and simplify the difference quotient of f, f(x+h) + f(x) / h, h is not equal to 0, for the function. f(x) = 4x + 9

jimthompson5910 (jim_thompson5910):

should be f(x+h) - f(x) / h right?

jimthompson5910 (jim_thompson5910):

ok, thx

OpenStudy (anonymous):

No, thank you

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = 4x + 9\] \[\Large f(x+h) = 4(x+h) + 9\] \[\Large f(x+h) = 4x+4h + 9\] -------------------------------- \[\Large \frac{f(x+h)-f(x)}{h}\] \[\Large \frac{(4x+4h + 9)-(4x+9)}{h}\] \[\Large \frac{4x+4h + 9-4x-9}{h}\] \[\Large \frac{(4x-4x)+4h + (9-9)}{h}\] \[\Large \frac{\cancel{(4x-4x)}+4h + \cancel{(9-9)}}{h}\] \[\Large \frac{4h}{h}\] \[\Large \frac{4\cancel{h}}{\cancel{h}}\] \[\Large 4\] So \[\Large \frac{f(x+h)-f(x)}{h}=4 \ \text{where} f(x) = 4x+9\]

jimthompson5910 (jim_thompson5910):

So answer is choice C

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

np